1.    A man drops a ball off a cliff. If it is travelling 40m/s just before it hits the ground, how tall is the cliff? (a=9.81m/s^2).

 


1
Expert's answer
2021-12-09T16:49:02-0500
"v^2=2gh\\40^2=2(9.8)h\\h=81.6\ m"